### v1

@@ 4.5 Example fields
We illustrate the detected boxes from our reference model over a few example fields from the SDC1 image outside the training area in Fig. 20. We use the same field size as our training input window, namely $256{\times}256$ pixels. The first field represents the central region where our detector reaches peak detection performance. This field illustrates the capability of our detector to identify faint sources outside of its training interval (Sect. 4.2). A few sources are considered missed, but almost all of them are blended with sources that are already properly detected. The few false detections can be separated into three cases: i) it is near a properly detected source, but our NMS process did not remove it, ii) it is near a source, but the matching criterion failed, certainly due to the target flux being lower than the minimum accessible flux, or iii) it is near a compact signal that does not correspond to a source. The first and second cases are the most common.
@@
The second field illustrates the extreme case of a large resolved galaxy. As previously stated, there is no similar source in the training area that the network could have learned to identify. Therefore, missing the large source itself was expected. However, this field illustrates our detector capability to prevent false detections in such a context. There are only a few detections considered to be false, and most of them seem to represent a real underlying source, indicating that the matching criteria likely rejected it because of the predicted flux or size. Several blended sources are also properly detected on the edges of the large source. We purposely saturated all the images at the lower half of the input dynamic to improve the readability of the figure for faint sources. This explains why the detector manages to identify blended sources in bright regions that appear saturated. There are many missed sources in this context, which are just completely hidden by the bright source and, therefore, undetectable.
@@
The last field illustrates a region far from the image center but still in a primary beam regime where several sources can be detected. This is typical of the intermediate regime between the default training area and our noise-only areas. With this field, we can verify the capability of our detector to interpolate between the two training contexts. While most sources are properly detected, we observe that the number of false positives is slightly higher than in the first field while still preserving a relatively high purity. There are only a few considered missed, and their origin is likely the same as in the two other fields. This field is typical of the regions of the full field for which we identify an overdensity of false detection in Fig 21.
@@
Figure 21: Distribution of various kinds of ”objects” over the full SDC1 field, minus the training area when necessary, for the YOLO-CIANNA-ref model. The match and false detections are based on the scorer, while the missed ones are based on our selection function. This figure represents absolute numbers of detections and not relative ones, so the distribution is relevant, but the actual value is scaled by the total number of detections. The central coordinates of the field are RA=0 deg, Dec=-30 deg.
4.6 Full field distribution
@@
Using the match criterion from the scorer, we can look at the match distribution over the full field and the distribution of false positives and missed sources, which are all illustrated in Fig. 21. As expected, all distributions follow the primary beam imprint over the field. The striking feature of this figure is that the false positives seem more concentrated in a ring that spans from the farthest point of the training area up to the edge of the primary beam. This is a direct consequence of the training area definition, and this specific distribution is more pronounced when we do not include the two additional noise-only regions in our training sample (Sect. 3.5). The distribution of missed sources based on our selection function mostly follows the primary beam footprint, illustrating that our detector completeness is mostly uniform over the full field. Both the distribution of false and missed exhibit compact over-densities of a few pixels, corresponding to the position of the few very large and bright sources.
@@ 5.1 Challenge definition and scorer biases
The SDC1 successfully gathered research teams to adapt, develop, and evaluate detection methods on a controlled dataset with a fixed evaluation metric. Even now that the challenge itself has ended, it remains a very valuable dataset in its challenge form and even more valuable with the added information about the full True catalog. However, as we stated in Sect. 3.2, the use of 4GB images does not really prepare the community for handling data-intensive workloads. The SDC2 and SDC3 editions are more challenging in this regard. In addition, our study highlighted a few limitations that are the result of some design choices of the challenge. In this section, we discuss these specific aspects and also propose a few directions that could help the design of future challenges.
@@ 5.1.1 Training area
In Bonaldi et al. (2021), the authors highlight that the source classification aspect is made challenging due to the small number of detectable sources for which all the frequencies are available and the strong class imbalance. As the frequency increases, the size of the training area is also reduced, leading to a very small training sample for supervised learning methods. More generally, they do not illustrate or discuss the distribution of the true sources and false positives over the full image field to highlight the potential effects of the training area selection. We have extensively discussed how having only access to a central part of the image for training affects the generalization capability of our detector Sect. 3.5 and proposed a workaround by adding the two extra noise-only training regions. In Appendix. C, we explore an alternative training region that spans over a full field radius to encapsulate all the sensitivity regimes from the primary beam. We demonstrate that this simple change solved most of the related issues. Overall, the SDC1 falls in a classical definition hiccup regarding supervised learning approaches, that is, having a training sample that does not represent the final task to perform. Interestingly, this aspect was corrected for the SDC2 (Hartley et al., 2023) by providing the training example in the form of an independent data cube, which is another statistical realization of the same underlying simulation.
@@ 5.1.2 True catalog properties
We discussed in Sect. 3.4 how much the construction of the training sample matters, especially regarding the selection of detectable sources only. While sampling the transition between detectable and non-detectable sources with a sufficient margin is a good thing, keeping the vast majority of nondetectable objects in the reference True catalog can cause several issues. At testing time, it strongly increases the chance of random association. To counterbalance this effect, strong matching criteria have been used in the SDC1 scorer, which then causes the rejection of properly detected sources too often. We note that in Bonaldi et al. (2021), the authors indicate that background faint sources contribute to the realism of the image noise. Still, most of them could have been ignored during scoring to reduce the random matches over undetectable sources. In all cases, a selection function must be defined to select a specific point in the detectable to undetectable transition. In this regard, the provided True catalog could have contained more information directly from the simulation instead of just the source properties to predict. This would have allowed the construction of a more advanced selection function while not changing the amount of information in the image from which the detection has to be made.
Some other design choices do not directly impact the detector performances but could have been made differently to serve other purposes. For example, the True catalog contains the absolute real values for each source. An alternative approach would have been to provide a single noised realization for each target parameter and accompany them with uncertainty measurements. It would have improved the realism of the reference catalog, considering how target sources can be acquired from real observations, and it would have been possible to add uncertainty predictions to the task by including it in the scorer.
@@ 5.1.3 Scorer metric
The main impactful choices are naturally in the evaluation metric itself in the form of the scorer (Sect. 3.3). Firstly, several error thresholds are permissive. Consequently, multiple detection methods with different characterization capabilities will have similar scores for the corresponding source parameters. Secondly, several parameters use an asymmetrical relative error with a symmetrical score response function. This implies that underestimated predictions tend to achieve better scores. Finally, several error functions do not consider the structural limits of the predicted parameters, like resolution limits, which affect, for example, the $Bmaj$, $Bmin$, and $PA$ parameters. Reaching a minimum observational limit for a parameter while the target can be smaller results in a strong positive relative error that is more penalized. This is striking for the Flux, $Bmaj$, and $Bmin$ error distribution illustrated in Fig. 19. Because the score added by a properly detected source depends on its characterization score, successfully detecting difficult sources is poorly rewarded compared to detecting obvious ones. As we discussed in Sect. 4.4, since characterization scores are averaged over all the detected sources regardless of their intrinsic difficulty, a good detector with a good characterization capability will have a lower characterization score than a less complete detector with poor characterization capability. In this regard, one appropriate approach for comparing two method characterization capabilities would be to use subsets of sources identified by the two detectors. In this regard, we strongly encourage any future method application to the SDC1 dataset to include figures that represent the distributed parameter errors as a way to evaluate their characterization capability (e.g., Fig. 19).
@@
To maximize its score, a detector has to identify a maximum number of objects while preserving a given purity dictated by its average source characterization score. However, due to how the scorer was tuned, the accessible score range is quite narrow and does not change much with the actual characterization quality. It means that this challenge is, in fact, strongly detection-focused because an additional identified source always contributes a similar score while a false positive always contributes -1. It is equivalent to having set a fixed purity requirement and tasked teams to compete to achieve the best completion at this purity level. Still, the set of match criteria includes predicted characteristics like the flux and the size of the source. However, to prevent random association with the many non-detectable objects, the detection thresholds on these parameters have been set to strict values. As for the source characterization, the match criterion uses asymmetrical relative errors on quantities with instrumental limitations. When defining a match, the effect of the synthesized beam on the size and position has been accounted for, but nothing was done regarding the predicted flux. Consequently, sources that are perfectly detected regarding the instrument limits can still be considered false positives if their True catalog flux is slightly overestimated due to the relative error asymmetry. In addition, since these sources get lower characterization scores, the detection purity in this regime has to be higher to add a positive score, biasing the detector to remove them more than necessary. This results in underestimating the capability of very strong detectors in low SNR regimes, for which it would be very interesting to compare the various method capabilities.
@@
Another drawback of this scoring approach is that it expects a list of considered true detections, which does not exploit the capability of some methods to order all their detections by confidence. In classical computer vision challenges, they request a per-detection score that is sorted to construct a completeness-purity curve based on a match criterion. The integral of this curve is then used to evaluate the detector capability over its full sensitivity range. We provide more details on constructing the classical mean average precision (mAP) metric in Appendix. A.8.2. Interestingly, since the SDC1 scorer represents a matching criterion, it can be used to construct an AP metric. The match information can then be used to sample the completeness-purity curve. We present the raw and smoothed completeness-purity curve for our YOLO-CIANNA-ref model in Fig. 22, along with the binned purity as a function of the objectness for each detection unit. Since most of the objects in the simulation are non-detectable, the resulting AP is low, but it should remain sensitive to variations in detection capability. As a reference, we achieve ${\rm AP_{SDC1}}=14.3428\%$. While this is an interesting alternative detection metric, it does not account for the source characterization of the detected sources, except as a part of the match criteria.
Figure 22: Detector sensitivity curves for the YOLO-CIANNA-ref model. The top frame represents the raw and smoothed purity-completeness curve on which the AP is computed. The bottom frame represents the binned purity as a function of the objectness for all detection units. The bin size is 0.01, 0.05, and 0.1 for the small, intermediate, and large size-prior regimes, respectively.
@@ 5.2 Method weaknesses and possible improvements
Our YOLO-CIANNA method has been proven capable of reaching a very high score on the SDC1 challenge dataset. However, there are a lot of aspects that can be discussed and improved.
@@ 5.2.1 SDC1 specific limitations
In this study, we tackled only a sub-problem compared to the original challenge description. We only applied our method to the 560 MHz - 1000h image and excluded the classification and core fraction estimate. We could have used a combination of images at different frequencies as independent input channels, but we would have been limited to the overlapping area of these images. The combination of the 560 MHz and 1400 MHz images would be an interesting case study for our method as the highest frequency would help reduce source confusion and even allow a simple attempt of classification. Still, our application to the single 560 MHz - 1000h image was sufficient to demonstrate the capabilities of our YOLO-CIANNA source detector in a context with a high density of small sources and occasional blending. It also allowed us to illustrate the high source characterization accuracy of our method. Regarding the specific case of very large and bright sources that are not detected, we discussed that it results from the lack of similar examples in the training area content and that it is not an intrinsic limitation of our method. One approach to recover these objects in the current setup would have been to downsample the image to reduce the apparent size of all objects and apply the same trained detector to recover the large sources. As discussed in Sect. 3.8 and Appendix A.1, changing the apparent resolution of the objects has a significant impact on what is considered to be the instrumental limits, hence the absence of rescaling augmentation in our training. However, using rescaled input with a trained network to recover only sources that are far from the resolution limit is a viable solution. Regarding the predicted parameters, simple scaling rules could be used to reconstruct the original quantities. We did not include this approach in this study as it would increase the complexity of the prediction for only a marginal number of recoverable sources.
We exposed in Sect. 5.1.2 that adding noise and uncertainty to the parameters to predict in the training catalog would have been more realistic. From a pure regularization point of view, we tried adding random noise realizations to the target parameters in our augmentation during the training process. While it stabilizes the training, it slightly reduces the best achievable score. Still, it demonstrates that including the uncertainty range is technically straightforward. To properly use target uncertainty and include it in the scoring procedure, our method would have to predict uncertainty as well. By having dropout in our network architecture, our YOLO-CIANNA method is already capable of using MC-dropout to predict posterior distributions for all the predicted parameters, which scales with the uncertainty (Srivastava et al., 2014; Gal & Ghahramani, 2015).

### v2

@@ 4.5 Example fields
We illustrate the detected boxes from our model over a few example fields outside the training area in Fig. 20. We use the same field size as our training input window, namely $256{\times}256$ pixels. The first field represents the central high-sensitivity region where our detector reaches peak detection performance. This field illustrates the capability of our detector to identify faint sources outside its training interval (Sect. 4.2). A few sources are considered missed, but most of them are blended with sources already correctly detected. The few false detections can be separated into three cases: i) it is near a properly detected source, but our NMS process did not remove it, ii) it is near a source, but the matching criterion failed, certainly due to the target flux being lower than the minimum accessible flux, or iii) it is near a compact signal that does not correspond to a source. The first and second cases are the most common.
@@
The second field illustrates the extreme case of an extended resolved galaxy. As previously stated, there is no similar source in the training area that the detector could have learned to identify. Therefore, the large source itself was expected to be missed. However, this field illustrates our capability to prevent false detections in such a context. Only a few detections are considered false, and most of them seem to represent a real underlying source, indicating that the matching criteria likely rejected it because of the predicted flux or size. Several blended sources are also properly detected on the edges of the extended source. The detector manages to identify blended sources in bright regions of the extended central source that appear saturated. The background images in this figure are, in fact, oversaturated to increase contrast over faint sources. Using the actual input image provided to the network, these detected sources are also obvious to the eye. There are many missed sources in this context, which are completely hidden by the bright source and are therefore undetectable. These missed undetectable sources highlight a typical possible improvement of our training selection function, which could be modified in order to remove them.
@@
The last field illustrates a region far from the image center but still in a primary beam regime where several sources can be detected. This is typical of the intermediate regime between the default training area and our noise-only areas. With this field, we can verify the capability of our detector to interpolate between the two training contexts. We observe that most sources are properly detected. Only a few are considered missed, and their origin is likely the same as in the two other example fields. This field is typical of the regions of the whole image for which we identify a small overdensity of false detection in Fig 21.
@@
Figure 21: Two-dimensional histograms of various kinds of “objects” representing their distribution over the full SDC1 field for the YOLO-CIANNA model. The training area is masked when necessary. All densities are binned using the same 200x200 grid. The match and false detections are based on the scorer, while the missed ones are based on our selection function. The central coordinates of the field are $\textrm{RA}=0$ deg, $\textrm{Dec}=-30$ deg.
4.6 Whole image field distribution
@@
Using the match criterion from the scorer, we can look at the match distribution over the whole field and the distribution of false positives and missed sources, which are all illustrated in Fig. 21. All distributions follow the primary beam imprint over the whole field. The striking feature of this figure is that the false positives appear more concentrated in a ring that spans from the farthest point of the training area up to the edge of the primary beam. This is a direct consequence of the training area definition, and this specific distribution is more pronounced when we do not include the two additional noise-only regions in our training sample (Sect. 3.5). The distribution of missed sources based on our selection function mostly follows the primary beam footprint, illustrating that our detector completeness is uniform over the whole image field. The distribution of false and missed both exhibit compact over-densities of a few pixels, corresponding to the position of the few extended and bright sources.
@@ 5.1 Challenge definition and scorer biases
The SDC1 successfully gathered research teams to develop, adapt, and evaluate detection methods on a controlled dataset with a fixed evaluation metric. Even now that the challenge has ended, it remains a valuable dataset in its challenge form, even more so now that they have provided the full True catalog. However, as stated in Sect. 3.2, with only 4GB images, the SDC1 does not prepare the community for handling data-intensive workloads. The following SDC2 and SDC3 editions are more challenging in this regard. In addition, our study highlighted a few limitations that are the result of some design choices of the challenge. In this section, we discuss these limits and propose a few directions that could help design future challenges.
@@ 5.1.1 Training area
In Bonaldi et al. (2021), the authors highlight that the source classification aspect is challenging due to the small number of detectable sources for which all frequencies are available. As the frequency increases, the size of the training area is also reduced, leading to a very small training sample, especially for rare classes. More generally, they do not illustrate or discuss the distribution of the matched sources and false positives of the submitted catalogs over the whole image field to highlight the potential effects of the training area selection. We extensively discussed how having only access to a central part of the image for training affects the generalization capability of our detector Sect. 3.5 and proposed a workaround by adding the two extra noise-only training regions. In Appendix. C, we explore an alternative training region that spans an entire beam radius to encapsulate all the sensitivity regimes from the primary beam. We demonstrate that this simple change solved most of the related issues. Overall, the SDC1 falls under the classical supervised learning hiccup of having a training sample that does not represent the final task to perform. Interestingly, this aspect was corrected for the SDC2 (Hartley et al., 2023), for which the training sample was provided as an independent data cube representing another realization of the same underlying simulation.
@@ 5.1.2 True catalog properties
In Sect. 3.4, we discussed how the construction of the training sample matters, especially regarding the selection of detectable sources. While sampling the transition between detectable and nondetectable sources with a sufficient margin is a good thing, keeping the vast majority of un-detectable objects in the reference True catalog causes several issues. At testing time, it increases the chance of random association. To counterbalance this effect, restrictive matching criteria have been used in the SDC1 scorer, notably on the flux, which then causes the rejection of actually detected sources too often (Sect. 5.1.3). We note that in Bonaldi et al. (2021), the authors indicate that background faint sources contribute to the realism of the image noise. Still, most of them could have been ignored during scoring to reduce the random matches over undetectable sources. In all cases, a selection function must be defined to select a fixed point in the detectable to undetectable transition. In this regard, the provided True catalog could have contained more information directly from the simulation instead of just the source properties to predict. For example, the underlying simulation produces pixel masks for all sources, which would help recognize blending or identify compact detectable regions in extended resolved sources. Such extra information provided only for the True catalog would have allowed the construction of a more advanced selection function while not changing the amount of information in the image from which the detection has to be made.
Some other design choices do not directly impact the detector performances but could have been made differently to serve different purposes. For example, the True catalog contains a unique true value for all the parameters of every source. An alternative approach would have been to provide a single noised realization for each target parameter and accompany them with uncertainty measurements. It would have improved the realism of the produced catalog, considering how target sources can be acquired from real observations. Consequently, uncertainty predictions could have been added to the scorer.
@@ 5.1.3 Scorer metric
The main impactful choices are naturally in the evaluation metric itself in the form of the scorer (Sect. 3.3). Firstly, several error thresholds are permissive. Therefore, multiple detection methods with different characterization capabilities will have similar scores for the corresponding source parameters. Secondly, most parameters use an asymmetrical relative error with a symmetrical score response function. This implies that underestimated predictions tend to achieve better scores. Finally, the error functions do not consider some structural or observational limits of the predicted parameters. For example, the limited instrumental resolution constrains the accessible values for the Bmaj, Bmin, and PA parameters, while the noise limit constrains the accessible flux values. Having target values lower than the observational limits results in an uncapped positive relative error that does not reflect the intrinsic detector capability. This is striking for the Flux, Bmaj, and Bmin error distribution illustrated in Fig. 19. Since the score of a properly detected source depends on its characterization, successfully detecting difficult sources is poorly rewarded compared to detecting obvious ones. Another issue specific to the flux is that the target value is the intrinsic source flux instead of the apparent flux. As the sensitivity varies with the primary beam, the flux score estimate and matching criteria are inhomogeneous over the image. This can also contribute to explaining the faint outer ring of false detection visible in Fig. 21. As discussed in Sect. 4.4, characterization scores are averaged over all the detected sources regardless of their intrinsic difficulty. Consequently, a good detector with a good characterization capability will have a lower average characterization score than a less complete detector with poor characterization capability. A better approach would be to use subsets of sources identified by the two detectors. In this regard, we strongly encourage any future method application to the SDC1 dataset to include figures that represent the distributed parameter errors as a way to evaluate their characterization capability (e.g., Fig. 19).
@@
To maximize its score, a detector has to identify as many objects as possible while preserving a given purity dictated by its average source characterization score. This challenge is strongly detection-focused as all identified sources always contribute a similar score, while a false positive always contributes -1. It is equivalent to having set a fixed purity requirement and tasked teams to achieve the best completion at this purity level. Still, the set of match criteria includes predicted characteristics like the flux and the size of the sources, which are used to prevent random association with the many nondetectable objects. As for the source characterization, the match criterion uses asymmetrical relative errors on quantities with instrumental limitations. When defining a match, the effect of the synthesized beam on the size and position has been accounted for, but nothing was done regarding the predicted flux. Consequently, perfectly detected faint sources regarding the instrument limits can be considered false positives if their True catalog flux is slightly overestimated. In addition, since these sources get lower characterization scores, the detection purity in this regime has to be higher to add a positive score, biasing the detector to remove them more than necessary. This results in underestimating the capability of strong detectors for low S/N, which is exactly the regime in which the comparison of method capabilities would be interesting. While using the flux in the matching criteria is a good idea to reduce the chance of random matches, it would have been better to incorporate the matching score in the global scoring instead of rejecting sources based on an arbitrary threshold. Also, the target apparent flux correlates well with the detection difficulty and could have been used as a conditional threshold for the match criteria. Another approach could have been to directly scale the contribution of each target to the total score based on their target flux.
5.1.4 Alternative custom AP metrics
SDC1 scorer matching
DIoU-based matching
Figure 22: Detector sensitivity curves for the YOLO-CIANNA model. The top frames represent purity-completeness curves built from running purity and precision scores on an objectness-sorted prediction list. The smoothed version considers that the purity at any completeness point is the maximum purity at any superior completeness. The bottom frames represent the binned purity as a function of the objectness for each detection unit. The bin size is 0.01, 0.05, and 0.1 for the small, intermediate, and large size-prior regimes, respectively. The left and right columns present curves produced with the SDC1 scorer matching rule and a purely geometrical DIoU-based with a 0.1 threshold matching rule, respectively. All curves are produced from the same candidate catalog where all detections with an objectness less than 0.03 have been removed.
@@ 5.1.4 Alternative custom AP metrics
Another drawback of this scoring approach is that it expects a list of considered true detections, which does not exploit the capability of some methods to produce a continuous confidence score. In classical computer vision challenges, the predicted score is used to construct a completeness-purity curve based on a relatively permissive match criterion. The integral of this curve is then used to evaluate the detector capability over its full sensitivity range (Appendix. A.8.2). For some challenges, they also build a metric that is the average of sensitivity curve integrals using multiple matching thresholds. Interestingly, since the SDC1 scorer represents a matching criterion, it can be used to sample the completeness-purity curve and build a custom AP metric. We present the raw and smoothed completeness-purity curve for our YOLO-CIANNA model over the test region in Fig. 22, along with the binned purity as a function of the objectness for each detection unit. To speed up the scorer cross-match, we rejected all predicted sources with an objectness score below $0.03$, which cuts the tail of the curve in the high-recall law precision regime, slightly reducing the AP score as well. We see that the purity drops quickly to a $0.9$ plateau up to $0.1$ completeness. This indicates that some confidently detected sources are rejected by the SDC1 scorer match criteria, which is likely the result of the scorer limitations discussed in Sect. 5.1.3. As a reference, we achieve ${\rm AP_{SDC1}}=14.47$. Since most of the objects in the simulation are nondetectable, the resulting AP is low, but it should remain sensitive to variations in detection capability. While this is an interesting alternative detection metric, it only accounts for the characterization of the sources through the match criteria.
While the source characterization is a core part of the SDC1 challenge and is likely to be required as well for observational data, we can try to evaluate how our method performs with a purely geometric detection-only metric. For this, we use the classical AP definition with a DIoU as our matching criteria. This allows us to focus mostly on source positioning while still accounting for the box size to some extent. Since our target boxes are small, this metric should only be lightly affected by random matches. We build the completeness-purity curve over the full image, which should not cause issues as the training area represents only a small fraction of the image, and we do not observe any sign of over-training. We also rejected all predicted sources with an objectness score below $0.03$. We present the sensitivity curve of this alternative metric in Fig. 22, along with the binned purity as a function of the objectness for each detection unit. As a reference, we achieve ${\rm AP_{DIoU:0.1}}=17.87$. We note that a DIoU limit of $0.1$ is already quite demanding regarding position accuracy, especially for small objects, as illustrated in Fig. 25. We see that the purity is near the maximum value up to a completeness of $0.1$, showing that at this DIoU threshold, all confident detections are considered to match a real source. However, for higher completeness values, the purity drops significantly, indicating that the low objectness source remains less reliable.
At this point, one can wonder if our choice of merging the source detection and characterization in a single network model was the best approach. As detection-related characteristics and extra-parameters contribute to a common loss (eq. 14), it would be possible that the network struggles to find a balance, reducing its predictive capability for both. However, in our experience, trying to optimize correlated parameters simultaneously usually produces better results than when optimizing them independently. As stated before, we consider the flux and source size to correlate with source detectability. Thanks to the detection-only AP metric we just defined, we can verify that a network trained to do the detection and characterization simultaneously performs better in terms of detection-only performances than a network trained on the detection task only. This demonstration is made in Appendix D.
Finally, we note that the obtained detection-only AP cannot be directly compared with the SDC1 equivalent. However, it shows that our method produces a high-quality detection and allows for future comparisons with other methods. For now, it is impossible to compute this metric using catalogs from other teams listed in Sect. 4.2, as the SDC1 file format does not ask for the required per-source detection score. These catalogs are also cut to optimize the SDC1 score and therefore do not sample the full sensitivity range of the associated detectors. We do not provide our full sensitivity catalog as an end product, but it can easily be reconstructed from the provided trained network models and associated scripts.
@@ 5.2.1 SDC1 specific limitations
In this study, we tackled only a subproblem compared to the original challenge description. We applied our method to the single 560 MHz - 1000h image and excluded the classification and core fraction estimate. We could have used a combination of images at different frequencies as independent input channels, but we would have been limited to their overlapping area. The combination of the 560 MHz and 1400 MHz images would be an interesting case study for our method as the highest frequency would help reduce source confusion and even allow a simple attempt of classification. Still, our application to the single 560 MHz - 1000h image was sufficient to demonstrate the capabilities of our YOLO-CIANNA source-detector in a context with a high density of small sources and occasional blending. It also demonstrated the characterization capability of our method.
Regarding the specific case of very extended and bright sources that are not detected, we discussed that it results from the lack of similar examples in the training area content and is not an intrinsic limitation of our method. One approach to recover these objects in the current setup would have been to downsample the image to reduce the apparent size of all objects and apply the same trained detector to recover the extended sources. As discussed in Sect. 3.8 and Appendix A.1, changing the apparent resolution of the objects has a significant impact on what is considered to be the instrumental limits, hence the absence of rescaling augmentation in our training. However, using rescaled input with a detector trained to recover only sources that are far from the resolution limit is a viable solution. Regarding the predicted parameters, simple scaling rules could be used to reconstruct the original quantities. We did not include this approach in this study as it would increase the complexity of the prediction for only a marginal number of recoverable sources.
We exposed in Sect. 5.1.2 that adding noise and uncertainty to the parameters to predict in the training catalog would have been more realistic. Following this idea, we tried adding random noise realizations to the target parameters in our augmentation during the training process as a regularization. While it stabilizes the training, it also reduces the best achievable score. Still, it demonstrates that accounting for the uncertainty is technically straightforward. If predicted uncertainty were included in the scoring procedure, our method would have to predict uncertainty as well. By adding dropout in our network, our YOLO-CIANNA method can already use MC-dropout to predict posterior distributions for all the predicted parameters, which scales with the uncertainty (Srivastava et al., 2014; Gal & Ghahramani, 2016).